OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of . Explain in your own words how you can simplify without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

 equaion

OpenStudy (anonymous):

x^2 ?

OpenStudy (btaylor):

1: \[\frac{(x)(x)(x)(x)}{(x)(x)}\rightarrow (x)(x) \rightarrow x^2\]Without expanding, just subtract the exponents. 2:\[\frac{2y^3}{4y^{-2}}\rightarrow \frac{y^5}{2}\]